# Efficient and Reliable Hitting-Set Computations for the Implicit Hitting Set Approach

### v1

@@ 0.4.2 Proofs and IP Instantiations of Solve-HS
The first approach invokes a proof-producing optimizer whenever the lower bound would be increased to equal the upper bound, i.e., when an optimal solution to the overall instance has been found. Intuitively, this represents the minimum amount of exact computation in Solve-HS needed to obtain a proof of correctness for the overall IHS algorithm; however, no guarantees on the intermediate lower bounds are obtained. Additionally, reinvoking a proof-producing optimizer when the LB would be increased might require some amount of duplicate work in the optimizer recomputing a solution already found by the IP or SLS solver.
@@
The second approach uses a proof-producing optimizer whenever the lower bound would be refined, i.e whenever the if statement on Line 2 of Algorithm 2 is true. Verifying all of the lower bounds requires more (potentially duplicated) exact computation. Nevertheless, this allows for generating proofs for more search heuristics used by instantiations of IHS that we are aware of.
@@
The previous two hybridizations possibly perform duplicate work (asking a proof-producing optimizer to confirm all/some optimality claims of other optimizers), which might be undesirable. In this third hybridization, we make an a priori distinction between calls where optimality is needed and calls where this is not needed. More specifically, this version uses proof-producing optimizers to only instantiate the calls to Optimize in which the parameter opt? is true. The calls to Optimize that are not required to find an optimal solution can be instantiated with non-proof producing optimizers. This ensures that any calls to Solve-HS that are required to find an optimal solution are indeed proof-producing. In this setting, the case $\gamma(O)=ub$ of the if statement on Line 2 that represents Optimize returning an optimal solution even when opt? is false, is not used since non-proof-producing optimizers may make such calls.
@@ 0.5 Experiments
In our experiments, we evaluate the performance of the different instantiations of Optimize as well as the overhead caused by proof logging. First, we compare how different solvers used as the single solver for Optimize perform as well as the hybrid configurations outlined in Section 0.4. Secondly, we evaluate the benefits of integrating SLS. Thirdly, we investigate the impact of proof logging on runtime performance. All implementation code will be made available in open source, together with run scripts and raw data.
@@
We run our experiments on compute nodes with two AMD EPYC 7513 CPUs, each with 64 cores running at 2.6 GHz. We reduce runtime variance by reserving whole CPUs for our experiments and run 16 instances in parallel on one CPU. Each run is limited to 14 GB memory and 1 h runtime using runlim [^3] https://github.com/arminbiere/runlim, and each solver is executed once on each instance.
@@
We use Gurobi 11.0.3 and SCIP 9.2.3444https://github.com/scipopt/scip (commit 3045f20) [BBB+24] as ILP solvers. The latter offers an exact mode (see Section 0.3.1). Further, we use a modified version of Roundingsat [^5] https://gitlab.com/MIAOresearch/software/roundingsat (Commit: c548e109) [GN21, BGMN23] as the proof producing PB solver and VeriPB 2.2.2666https://gitlab.com/MIAOresearch/software/VeriPB to check these proofs. Proof logging is enabled only for experiments concerning proof logging.
@@ 0.5.1 Impact of Hitting Set Oracles
Table 1: Different instantiations of Optimize. Single shows instantiations using only a single solver and Hybrid shows instantations that combine an ILP solver (Gurobi) with a proof producing solver (Roundingsat). #: number of instances solved, m: average memory usage in GB.
In our first experiment, we use a single solver for the hitting set computation in Optimize. Table 1 shows the results for different solvers, split into ILP solvers (top) and designated PB solving methods. These results show that the ILP solvers solver more instances. However, the difference between SCIP and Exact SCIP show that guaranteed optimality causes a significant performance decrease. Indeed, the number of solved instances decreases by more than 15% while the memory usage doubles.
The dedicated PBO methods solve fewer instances, motivating the hybrid approaches. The further we delay the call to RoundingSat, the better the results: AllLB has the most overhead and performs overall poorest, while OptLB requires the fewest RoundingSat calls and performs overall best. Surprisingly, SIS with core boosting does not perform well on its own, being the second-worst method on its own, but performs best in our hybrid approaches. The difference might be due to an overhead in core boosting that only affects performance in frequent incremental calls.
@@ 0.5.2 Impact of Integrating Local Search
We integrated Core-SLS (cf. Section 0.3.2) into the single solver instantiations of Optimize and the best hybrid approach OptLB. By Table 2, SLS improves all instantiations. We use two SLS configurations, one for SLS combined with an ILP solver and one for SLS combined with dedicated PBO approaches. In the hybrid configuration, we use the SLS configuration for ILP. Integration of SLS consistently results in more instances solved, decrease memory usage and improve runtimes. This is most striking for Exact SCIP, for which SLS gives the biggest gain in the number of solved instances, also decreasing the memory consumption by almost 20%. Interestingly, often large numbers of instances are solved either with or without SLS but not both (Column d in Table 2). However, when SLS does not find a sufficiently low-cost solution it directly incurs a runtime overhead. Despite diversification, SLS can get stuck in local optima and cannot escape it without an optimizer. This suggests that improving SLS heuristics further for the IHS use case has potential for even further improvements.
In the hybrids, SLS can only replace ILP solver calls due to inexactness. Here, SLS contributions significantly decrease, as the number of instances that SLS performs well on is balanced with instances it does poorly on. It is unlikely that instances the ILP solver can only solve when SLS is used could be solved by the dedicated PBO approach.
@@ 0.5.3 Proof Logging Overhead
We evaluate proof logging using the most performant proof-producing variant, OptLB using SIS + CB. The proofs are checked using VeriPB with a timelimit of 10 h. By proof logging, we found several occasions where the ILP solver wrongly claimed optimality; Gurobi on four instances a total of nine times, and four wrong results for Exact SCIP.

### v2

@@ 0.4.2 Proofs and IP Instantiations of Solve-HS
The first approach invokes a proof-producing optimizer whenever the lower bound would be increased to equal the upper bound, i.e., when an optimal solution to the overall instance has been found. Intuitively, this represents the minimum amount of exact computation in Solve-HS needed to obtain a proof of correctness for the overall IHS algorithm. However, no guarantees on the intermediate lower bounds are obtained. Additionally, re-invoking a proof-producing optimizer when the LB would be increased might require some amount of duplicate work in the optimizer recomputing a solution already found by the IP or SLS solver.
@@
The second approach uses a proof-producing optimizer whenever the lower bound would be refined, i.e., whenever the if statement on Line 2 of Algorithm 2 is true. Verifying all of the lower bounds requires more (potentially duplicated) exact computation. Nevertheless, this allows for generating proofs for more search heuristics used by instantiations of IHS that we are aware of.
@@
The previous two hybridizations possibly perform duplicate work (asking a proof-producing optimizer to confirm all/some optimality claims of other optimizers), which might be undesirable. In this third hybridization, we make an a priori distinction between calls for which optimality is required and calls for which optimality is not required. Specifically, this version uses proof-producing optimizers to only instantiate the calls to Optimize in which the parameter opt? is true. The calls to Optimize that are not required to find an optimal solution can be instantiated with non-proof producing optimizers. This ensures that any calls to Solve-HS that are required to find an optimal solution are indeed proof-producing. In this setting, the case $\gamma(O)=ub$ of the if statement on Line 2 that represents Optimize returning an optimal solution even when opt? is false, is not used since non-proof-producing optimizers may make such calls.
@@ 0.5 Experiments
In our experiments, we evaluate the performance of the different instantiations of Optimize as well as the overhead caused by proof logging. First, we compare how different solvers used as the single solver for Optimize perform as well as the hybrid configurations outlined in Section 0.4. Second, we evaluate the benefits of integrating SLS. Finally, we investigate the impact of proof logging on runtime performance. All implementation code, together with run scripts and raw data, can be found online [IVS+25].
@@
The experiments reported on were run on compute nodes with two AMD EPYC 7513 CPUs, each with 64 cores running at 2.6 GHz. We reduced runtime variance by reserving whole CPUs for our experiments and run 16 instances in parallel on one CPU. Each run is limited to 14 GB memory and 1 h runtime using runlim [^2] https://github.com/arminbiere/runlim, and each solver is executed once on each instance.
@@
All our implementation builds on the PBHS solver. [^3] https://bitbucket.org/coreo-group/pbhs We use Gurobi 11.0.3 and SCIP 9.2.3444https://github.com/scipopt/scip (commit 3045f20) [BBB+24] as IP solvers. The latter offers an exact mode (see Section 0.3.1). Further, we use a modified version of Roundingsat [^5] https://gitlab.com/MIAOresearch/software/roundingsat (Commit: c548e109) [GN21, BGMN23] as the proof producing PB solver and VeriPB 2.2.2666https://gitlab.com/MIAOresearch/software/VeriPB to check these proofs. Proof logging is enabled only for experiments concerning proof logging.
@@ 0.5.1 Impact of Hitting Set Oracles
Table 1: Different instantiations of Optimize. Single shows instantiations using only a single solver and Hybrid shows instantations that combine an IP solver (Gurobi) with a proof producing solver (Roundingsat). #: number of instances solved, m: average memory usage in GB.
In the first experiment, we use a single solver for the hitting set computation in Optimize. Table 1 shows the results for different solvers, split into IP solvers (top) and designated PB solving methods. These results show that the IP solvers solve more instances. However, the difference between SCIP and Exact SCIP show that guaranteed optimality causes a significant performance decrease. Indeed, the number of solved instances decreases by more than 15% while the memory usage doubles.
As far as the dedicated PBO methods are concerned, first we notice that the variant of solution-improving search with reified solution-improving constraints appears to suffer from too much memory usage and is outperformed by pure SIS. This difference disappears when using the OptLB hybrid strategy and the number of calls to the PBO oracle is significantly decreased. Over the line, we observe that the combination of solution-improving search with core-boosting outperforms the SIS as well as the core-guided implementation.
The dedicated PBO methods solve fewer instances than the IP solvers, but they do come with certification, motivating the hybrid approaches. The more we delay the call to RoundingSat, the better the results: AllLB has the most overhead and performs overall poorest, while OptLB requires the fewest RoundingSat calls and performs best overall.
@@ 0.5.2 Impact of Integrating Local Search
Secondly, we integrated Core-SLS (cf. Section 0.3.2) into the single solver instantiations of Optimize and the best hybrid approach OptLB. By Table 2, integration of SLS improves the runtime performance of each of the instantiations. We use two SLS configurations, one for SLS combined with an IP solver and one for SLS combined with dedicated PBO approaches. In the hybrid configuration, we use the SLS configuration for IP. Integration of SLS consistently results in more instances solved, decrease memory usage and improve runtimes. This is most striking for Exact SCIP, for which SLS gives the biggest gain in the number of solved instances, also decreasing the memory consumption by almost 20%. Interestingly, often large numbers of instances are solved either with or without SLS but not both (Column d in Table 2). However, when SLS does not find a sufficiently low-cost solution it directly incurs a runtime overhead. Despite diversification, SLS can get stuck in local optima and cannot escape it without an optimizer. This suggests that improving SLS heuristics further for the IHS use case has potential for even further improvements.
In the hybrids, SLS can only replace IP solver calls due to inexactness. Here, SLS contributions significantly decrease, as the number of instances that SLS performs well on is balanced with instances it does poorly on. It is unlikely that instances the IP solver can only solve when SLS is used could be solved by the dedicated PBO approach.
@@ 0.5.3 Proof Logging Overhead
We evaluate proof logging using the most performant proof-producing variant, OptLB using SIS + CB. The proofs are checked using VeriPB with a timelimit of 10 h. By proof logging, we found several occasions where the IP solver wrongly claimed optimality; Gurobi on four instances a total of nine times, and four wrong results for Exact SCIP.
@@ 0.6 Conclusion
Acknowledgments
This work is partially funded by the European Union (ERC, CertiFOX, 101122653). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
In addition, this work is partially funded by the Fonds Wetenschappelijk Onderzoek – Vlaanderen (G064925N and G070521N & fellowship 11A5J25N) and by the Research Council of Finland (grants 362987 and 356046), and by an Amazon Research Award (Fall/2023).
The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG.